The length and breadth of a rectangle are both doubled. What effect does this have on the perimeter and area of the rectangle
Answer by Alan3354(69443)

The length and breadth of a rectangle are both doubled. What effect does this have on the perimeter and area of the rectangle
------------
Area = L*W
P = 2L + 2W = 2*(L+W)
------------------
Use 2L and 2W